What would be the effective rate of protection on bicycles in China if China places a 50 percent tariff on bicycles, which have a world price of Php200,and no tariff on bike components, which together have a world price of Php100?



1
Expert's answer
2021-03-29T07:21:36-0400

Solution:

First, calculate the value-added before tariffs and after tariffs:


Value-added (VA) = World price of output – Cost of inputs (components)


Value added without tariff:

Bicycle world price = Php200

Component world price = Php100

Value added = 200 – 100 = Php100


Value-added with tariff:

Bicycle world price = "200 + (200\times50\%) = 200+100 = Php300"

Component world price = Php100

Value added = 300 – 100 = Php200

Effective rate of protection of bicycles (g) ="\frac{(VA \;after\; tariffs - VA\; added\; before\; tariffs) }{VA \;added \;before\; tariffs} \times 100"


g = "\frac{(200 - 100) }{100} \times 100 = 100\%"


The effective rate of protection of bicycles = 100%